OpenStudy (anonymous):

Calculate the integral over the bounded region: ∫B sqrt(x^2+y^2) dxdy, B={(x, y) ∈ R^2 |0 ≤ x ≤ y, x^2+ y^2≤ 1}

OpenStudy (anonymous):

\[\int\limits_{}^{}B \sqrt{x^2+y^2} dx dy, B=\left\{ (x,y) \in R^2|0\le x \le y, x^2+y^2\le 1 \right\}\]

OpenStudy (anonymous):

any help will be appreciated.